# Re-examining learning linear functions in context

### v1

@@ 1 Introduction
In-context learning (ICL) Brown et al. (2020) promises to make interacting with LLMs easy and accessible. ICL enables the model to learn a task from a prompt with instructions and a few examples at inference time, without any adjustment of the model’s parameters from pretraining. While there have been theoretical reconstructions of ICL, there have been few studies on exactly how ICL works in practice. However, ICL depends on a model’s pretraining as Garg et al. (2022) have shown; so doing an in depth analysis of this feature of LLMs is difficult. Hence, most of analysis done on how ICL works are done on small models and simple tasks. Garg et al. (2022) makes the problem mathematically precise: the model learns a task/function given in-context examples at inference time in a next-token-prediction format Brown et al. (2020); given a prompt containing a task input-output examples $(x_{1},f(x_{1}),..,x_{n},?)$, the model is asked to generate a value approximating $f(x_{n})$.
@@
  1. 1.
    Several recent papers claim that Transformer based models trained from scratch can ICL linear functions with performances close to algorithms such as Least Squares, and that a transformer can implement this algorithm. We show that the models we tested do not do this; they also failed to generalize and to provide robust predictions beyond their training data. In particular, all our transformer models failed to ICL the concept of a strictly increasing or strictly decreasing linear function, even over larger intervals in $\mathbb{R}$. We replicated this behavior for several training and test distributions. We trained transformers on different distributions various Gaussian, Bimodal and Uniform distributions.
  2. 2.
    Our experiments show that all our models trained from scratch have ‘boundary values” ($B,-B$) for prompts $x_{i}$; when $f(x_{i})>B$ or $<-B$, model performance degrades substantially. Training on uniform distributions makes this particularly clear.
  3. 3.
    All our transformer models solve task of ICL linear function by learning a projection from “nearby” sequences of points in the training data; Section 5 provides a mathematical formulation of what we think the models are doing. The exact projection depends upon the training distribution. The models do not implement linear regression or an algorithm like linear interpolation.
@@ 2 Background
Statistical learning examines the application of a learned function over a test domain and the expected loss over novel applications. The ability to bring the error over test to that over the training set is typically taken to indicate an ability to generalize.
@@
The class of linear functions ${\cal L}$ is clearly uniformly learnable. What is left open here is the choice of distribution of the data both for train and test and the sampling method (since our class is uncountably large). Garg et al. (2022) take a definition of learning where average expected error goes to 0 when data in train and test are sampled both from the same normal distribution. So for instance this means that the model will see linear functions with $a,b\in[-1,1]$ around 70% of the time. This can make a big difference on the error reported for learning a class of mathematical functions like ${\cal L}$, whose definition does not in any way depend on a particular distribution or sampling. Nevertheless, we would hope the model has found an algorithm such that $\hat{f}=f$ given a test set of linear functions with $a,b\not\in[-1,1]$. We also hope the algorithm will transfer to different distributions. This is what we investigate below.
@@ 3 Related Work
Further research then offered several theoretical reconstructions for how ICL for linear functions might work in Transformers. Von Oswald et al. (2023); Ahn et al. (2023); Mahankali et al. (2023) provided a construction to show transformers ICL from their doing gradient descent during ICL. Fu et al. (2023) showed that Transformers could ICL in virtue of using higher-order optimization techniques. Xie et al. (2021); Wu et al. (2023); Zhang et al. (2023); Panwar et al. (2023) argued that ICL follows on a Bayesian point of view. Bai et al. (2024) show that transformers can under certain assumptions implement many algorithms with near-optimal predictive power on various in-context data distributions. Giannou et al. (2024); Zhang et al. (2024) modify transformers with linear attention and Zhang et al. (2024) introduce a new training regime to show that modified transformers can learn linear functions. Given Pérez et al. (2021)’s result that full transformers with linear attention are Turing complete, however, these theoretical demonstrations are perhaps not surprising. Xie et al. (2021); Zhang et al. (2024) examine how ICL works despite differences between training and inference distributions.
@@ 4.1 Models do not implement linear regression
The results in Figure 1 and Table 2 (Table 2 is in Appendix B) confirm that at least the larger models are able to generalize somewhat to unseen examples, given that all the curves in Figure 1 have lower error rates than the baseline that predicts $\hat{f}(x_{n})=0$ everywhere. But their generalizing ability was far from perfect; and contrary to what Akyürek et al. (2022); Von Oswald et al. (2023) have suggested, the models did not use linear regression to ICL the target function. If they had, we would not see the error patterns we do.
Our models’ performance depends on how often it has seen examples “similar” to the target function value it is trying to predict. When $D_{F}=N(0,\sigma)$ there is an over 68% chance that a function chosen for train $f(a,b)$ will have $a,b\in[-\sigma,\sigma]$ and over a 95% chance it will have $a,b\in[-2\sigma,2\sigma]$. So a model with $D_{F}=D_{I}=N(0,1)$ has seen sequences of values for $f(a,b)$ with $a,b\in[-2,2]$ more than 95% of the time. Given a pretraining with over a billion examples, models will have seen prompts for functions with $a,b\not\in[-2,2]$, just not many of them. As the models are tested with $D^{t}_{F}=N(0,\sigma)$ and so confronted with more sequences representing functions $f(a,b)$ for $a,b\not\in[-2,2]$, all the models do less and less well.
@@
We’ve just seen that when the distribution of training data follows a simple Gaussian $N(0,1)$, the models, for any number of layers and attention head, give good results when $D^{t}_{F}=D^{t}_{I}=N(0,1)$, but offer degraded performance as we test on distributions $N(0,\sigma)$ for larger $\sigma$. We now show that same sort of behavior with different training distributions but all tested on $N(0,\sigma)$.
@@
We tested how our models fared with a bimodal distribution of training data. Our strategy was to increase the values the model can see during training, by extending the distribution seen during training. We trained our GPT2 on the bimodal distribution $0.5N(-1,1)+0.5N(1,1)$. With this training, the model was more likely to see wider values during training and therefore work better on larger values.
Most of the models we tested had more robust performance with a bimodal distribution for $D_{F}=0.5N(-1,1)+0.5N(1,1)$ than they did with $D_{F}=N(0,1)$ at least with $D^{t}_{F}=D^{t}_{I}=N(0,\sigma)$ and $n\geq 6$. The best models had almost equally good performance on $D^{t}_{F}=N(0,\sigma)$ for $\sigma\leq 3$ and superior performance with $D^{t}_{F}=N(0,\sigma)$ for $\sigma\geq 3$, as can be seen from Table 1. For the values of the table, we took $D_{I}^{t}=U(-1,1)$, the uniform distribution over $[-1,1]$, but the results remain similar when taking $D_{I}^{t}=N(0,1)$. The fact that performance varies with the distribution should not happen, if the models were using gradient descent to compute linear regression in ICL.
@@
To have more control on the notion of maximum and minimum values the models saw, we next trained our models on uniform distributions. We illustrate with $U(-5,5)$. Given the observations of Section 4.1 concerning the errors our models made on functions with large coefficients, we wanted to study whether these errors arose because the models hadn’t encountered functions with such large coefficients in pretraining. By keeping $D_{F},D_{I}$ normal or bimodal, we can’t control “the largest value the model could see”, because it’s always possible that it could have generated a large value during training. By training on a uniform distribution, however, we know exactly what the smallest and largest values that the model could have seen in its training. For example, setting $D_{F},D_{I}$ to $U(-5,5)$, the largest value the model could have seen is $30=5*5+5$ and the smallest value it could have seen is $-30$. Most likely it saw values significantly $>-30$ and $<30$.
Training with $U(-5,5)$ gave good results for $D^{t}_{F}=D^{t}_{I}=U(-1,1)$. Models were able to find target functions with coefficients in [-1,1] from only 2 points (see leftmost plot of Figure 6 in Appendix C); and all our models work well when $D_{F},D_{I},D^{t}_{F},D^{t}_{I}$ use the same distribution. The models trained on a uniform distribution sometimes do even better than models trained on N(0,1) or a bimodal distribution–up to three times better for $D^{t}_{F}=D^{t}_{I}=N(0,9)$ as Table 1 shows. Learning was at times very efficient, requiring just two prompts, as in Figure 6 (Appendix B).
@@
Table 1: Comparison showing the evolution of squared errors for models trained on different distributions; index N: $D_{F}=N(0,1)$, B $D_{F}=0.5N(-1,1)+0.5N(1,1)$ and $D_{F}=U(-5,5)$. We show error rates for models prompted without and with the natural ordering on the prompts [sorted], for the large model size. $D^{t}_{i}=U(-1,1)$ and $D^{t}_{F}=N(0,\sigma)$
@@ 4.3 Error analysis, sigmoid approximations and boundary values
We wanted to look at what sorts of errors our models made for $f(a,b)\in{\cal L}$ for $a,b\not\in[-2,2]$– i.e. the values of $\hat{f}(x)$ outside the interval that includes the vast majority they have seen. Our models exhibit problematic behavior of 2 kinds. Even our best models, for $a,b\not\in[-2,2]$ but reasonably close, say in $[-9,9]$, predict $\hat{f}(x)$ to a sigmoid-like function with correct estimates for the target function within a certain interval. [^2] Giannou et al. (2024) tested a different type of model, with linear attention only, and showed similar behavior. Consider Figure 2 for an illustrative $f(x)=8x+9$.
Figure 2: Plots of predictions of the 12L8AH model trained on $N(0,1)$ and error evolution over number of prompts for $f(x)=8x+9$
While the left plot of Figure 2 shows that the model’s prediction $\hat{f}(x)$ diverges dramatically from $f(x)$ outside of a certain interval, the right plot of Figure 2 shows the model has learned something with ICL about the function from the prompt and approximates it at least within a certain interval.
@@
For equations with coefficients highly unlikely to be sampled in $N(0,1)$ (for example $f(x)=30x+30$ in Figure 3), however, the results are catastrophic. Figure 3 shows that the model’s prediction $\hat{f}$ doesn’t converge to any stable prediction with ICL. This happens across a range of models.
@@
The same pattern holds when $D_{F},D_{I}=U(-5,5)$. Consider again as an illustrative example the target function, $f(x)=9x$ for our largest trained model. The model approximates $f(x)$ well within a certain range, but it predicts $\hat{f}(x)$ to be a constant function for intervals outside of that range. As shown in the left plot in Figure 7 in Appendix D, $\hat{f}^{+}(v)\approx 30$ for values $v$ for which the ground truth target function $f$ is such that $30\leq f(v)$, and the model predicts an approximally constant function $\hat{f}^{-}(v)\approx-30$ for values $v$ on which $f(v)\leq-30$. We call values like -30 and 30 where the model starts to predict constant functions $\hat{f}(v)$ and $\hat{f}^{-}(v)$ boundary values. For models trained on $U(-5,5)$, 30 and -30, with $30=5*5+5$ and $-30=-5*5-5$, are respectively the biggest and smallest values the model could have seen during training. If such a model hasn’t seen a value above 30 or below -30, it won’t infer one.
This behavior just noted for our largest model remains true for other models $M$ values $\hat{f}(v)<B-\alpha$ or $\hat{f}(v)>B+\alpha$, where $\alpha$ is a constant determined by $M$. However for functions and data samples when the values of $f(x)$ in the prompt sequence are such that $f(x)>B+\alpha$ or $<-B-\alpha$—an example is $f(x)=40x+40$—the model starts to make a mess of things, assigning $\hat{f}(v)$ random values for $f(v)$ far away from $B$ (i.e $>B+\alpha$ or $<-B-\alpha$, as seen in the rightmost plot in Figure 7. The model loses its capacity to model the target function anywhere.
@@
All our models trained on $U(-5,5)$ estimate the target function more or less well for $x$ with $f(x)\in[-B,B]$ with boundary values $-B,B$; but once we are outside $[-B,B]$, the estimations become constant functions or chaotic. Different models trained on different uniform distributions give different boundary values; for instance see Figures 3 or 10 Appendix F.
Boundary values for models on normal and bimodal distributions will vary, as the largest values in the sequences the model has seen and the number of times it has seen those values will vary depending on the sampling. We have found that larger models will have slightly larger boundary values $|B|$ than smaller ones. For instance, Figure 4 shows the plots for the predictions of two models (12L8AH, and 6L4AH) with both $D_{F},D_{I}=N(0,1)$ for the equation $f(x)=10x$. The larger model has boundary values $\approx$ -13.7, 13.7, the smaller one boundary values $\approx$ -12, 12.
Figure 4: Plots for $f(x)=10x$ by a 12l8ah model and by a 6l4ah model.
@@
Constraints from boundary values hold regardless of model size (for plots see Appendix D and Figure 8) and also hold for attention only models (See Appendix E, Figure 9). Larger models trained on the same distribution and the same number of data are able to ICL ${\cal L}$ functions over a slightly larger number of intermediate values than smaller models, as Figure 1 suggests.

### v2

@@ 1 Introduction
In-context learning (ICL) Brown et al. (2020) promises to make interacting with LLMs easy and accessible. ICL enables the model to learn a task from a prompt with instructions and a few examples at inference time, without any adjustment of the model’s parameters from pretraining. While there have been theoretical reconstructions of ICL, there have been few studies on exactly how ICL works in practice. ICL depends on a model’s pretraining; so doing an in depth analysis of this feature of LLMs is difficult. Hence, most of analysis done on how ICL works are done on small models and simple tasks. Garg et al. (2022) makes the problem mathematically precise: the model learns a task/function given in-context examples at inference time in a next-token-prediction format Brown et al. (2020); given a prompt containing a task input-output examples $(x_{1},f(x_{1}),..,x_{n},?)$, the model is asked to generate a value approximating $f(x_{n})$.
@@
1. Several recent papers claim that Transformer based models trained from scratch can through ICL implement algorithms like linear and ridge regression or Newton’s method. By shifting sampling from different training and test distributions of both functions $f$ and values $x_{i}$, we show that the models we tested do not do this and fail to generalize or to provide robust predictions beyond their training data. In particular, all our transformer models failed to ICL the concept of a strictly increasing or strictly decreasing linear function, even over larger intervals in $\mathbb{R}$. We trained transformers on different distributions various Gaussian, Bimodal and Uniform distributions.
2. Our experiments show that all our models on all training distributions (though training with uniform distributions makes this particularly clear) have ‘boundary values” ($B,-B$) for prompts $x_{i}$; when $f(x_{i})>B$ or $<-B$, model performance degrades substantially. We argue boundary values are crucial to understanding ICL.
3. All our transformer models solve the task of ICL linear function by learning a projection from “nearby” sequences of points in the training data.
@@ 2 Background
The class of linear functions ${\cal L}$ is clearly uniformly learnable. What is left open here is the choice of distribution of the data both for train and test and the sampling method (since our class is uncountably large). Garg et al. (2022) take a definition of learning where average expected error goes to 0 when data in train and test are sampled both from the same normal distribution. However, a class of mathematical functions like ${\cal L}$ does not in any way depend on a particular distribution or sampling. And so we would expect that if the model has ICL ${\cal L}$, it has found an algorithm such that $\hat{f}=f$ given a test set of linear functions and points not in its training distribution. In such a case the model will ICL with different distributions. This is what we investigate below.
@@ 3 Related Work
Further research then offered several theoretical reconstructions for how ICL for linear functions might work in Transformers. Von Oswald et al. (2023); Ahn et al. (2023); Mahankali et al. (2023) provided a construction to show transformers ICL from their doing gradient descent during ICL. Fu et al. (2023) showed that Transformers could ICL in virtue of using higher-order optimization techniques. Xie et al. (2021); Wu et al. (2023); Zhang et al. (2023); Panwar et al. (2023) argued that ICL follows from Bayesian principles. Bai et al. (2024) show that transformers can under certain assumptions implement many algorithms with near-optimal predictive power on various in-context data distributions. Given Pérez et al. (2021)’s result that full transformers with linear attention are Turing complete, however, these theoretical demonstrations are perhaps not surprising.
Xie et al. (2021); Zhang et al. (2024) show that when we shift training and inference distributions ICL performance degrades. Thus, this work is closer to our own as is Giannou et al. (2024). However, Giannou et al. (2024); Zhang et al. (2024) make important modifications to transformer architectures Giannou et al. (2024); Zhang et al. (2024) work with linear attention, whereas we look at attention layers as they actually are used with softmax. In addition, Zhang et al. (2024) uses a new kind of optimization or training with gradients and a special fixed initial point. This means that their architecture and training are quite different from what normally happens with transformers; they are interested in getting a revised transformer-like model to learn linear functions, while we want to find out whether transformers as they actually are learn linear functions or something else. As we detail below, the results for the architectures of Zhang et al. (2024); Giannou et al. (2024) are quite different from those we have for actual transformers. In addition unlike either of these papers, we show that prompts that are too long induce chaotic behavior.
@@ 4.1 Models do not implement linear regression
Observation 1
Transformer models can in principle ICL any linear function f approximable over an interval [a, b] if we pick a suitable training = testing regime.
@@
The results in Figure 1 and Table 2 confirm that at least the larger models are able to generalize somewhat to unseen examples, given that all the curves in Figure 1 have lower error rates than the baseline that predicts $\hat{f}(x_{n})=0$ everywhere. But their generalizing ability was far from perfect; and contrary to what Akyürek et al. (2022); Von Oswald et al. (2023) have suggested, the models did not use linear regression to ICL the target function. If they had, we would not see the error patterns we do.
Observation 2
All models tested failed to generalize in ICL to seldom seen linear functions $f$ or to use linear regression to predict $f$.
Zhang et al. (2024) claim shifting the distribution sampled at inference of the functions does not affect their models. Our results differ from theirs; such a shift affects the results in an important way, where we take $N(0,1)=D_{F}$ ( but $D^{t}_{F}=N(0,\sigma)$ for $1\leq\sigma\leq 10$. Figure 1 clearly shows that for transformer models with soft attention, this task shift reduces performance dramatically.
@@
We’ve just examined the behavior of models on test sampling from $N(0,\sigma)$ for larger $\sigma$ when the distribution of training data follows a simple Gaussian $N(0,1)$. Our models, for any number of layers and attention head, have the same behavior when trained on different distributions but tested on $N(0,\sigma)$; they give good results when $D^{t}_{F}=D^{t}_{I}=N(0,1)$, but offer degraded performance when tested on $N(0,\sigma)$ for larger $\sigma$.
@@
We tested how our models fared with the bimodal distribution of training data, $0.5N(-1,1)+0.5N(1,1)$. This increased the values of $f(x)$ the model can see during training.
Most of the models we tested had more robust performance with a bimodal distribution for $D_{F}=0.5N(-1,1)+0.5N(1,1)$ than they did with $D_{F}=N(0,1)$ at least with $D^{t}_{F}=D^{t}_{I}=N(0,\sigma)$ and $n\geq 6$. The best models had almost equally good performance on $D^{t}_{F}=N(0,\sigma)$ for $\sigma\leq 3$ and superior performance with $D^{t}_{F}=N(0,\sigma)$ for $\sigma\geq 3$, as can be seen from Table 1. For the values of the table, we took $D_{I}^{t}=N(0,1)$.The fact that performance varies with the distribution should not happen, if the models were using gradient descent to compute linear regression in ICL.
@@
We next trained our models on uniform distributions, in particular $U(-5,5)$. This gives more control on the notion of maximum and minimum values the models see in training. Given the observations of Section 4.1 concerning the errors our models made on functions with large coefficients, we wanted to study whether these errors arose because the models hadn’t encountered functions with such large coefficients in pretraining. By keeping $D_{F},D_{I}$ normal or bimodal, we can’t control “the largest value the model could see”, because it’s always possible that it could have generated a large value during training. By training on a uniform distribution, however, we know exactly what the smallest and largest values that the model could have seen in its training. For example, setting $D_{F},D_{I}$ to $U(-5,5)$, the largest value the model could have seen is $30=5*5+5$ and the smallest value it could have seen is $-30$. Most likely it saw values significantly $>-30$ and $<30$.
Training with $U(-5,5)$ gave good results for $D^{t}_{F}=D^{t}_{I}=U(-1,1)$. Models were able to find target functions with coefficients in [-1,1] from only 2 points (see leftmost plot of Figure 9 in Appendix C); and all our models work well when $D_{F},D_{I},D^{t}_{F},D^{t}_{I}$ use the same distribution. Learning was at times very efficient, requiring just two prompts, as in Figure 9 (Appendix B).
Observation 3
(i). All models with $D_{F}=U(-1,1)=D_{I}$ had lower error than when $D_{F}=U(-5,5)=D_{I}$
(ii) Given that the density of training sequences is higher with $D_{F}=U(-1,1)=D_{I}$ than with $D_{F}=U(-5,5)=D_{I}$, the density of training sequences affects model performance.
(iii). All models with $D_{F}=U(-5,5)=D_{I}$ had error values least as good as models with $D_{F}=N(0,1)=D_{I}$: They sometimes did up to 3 times better (cf values for $D^{t}_{F}=D^{t}_{I}=N(0,9)$ in Table 1)
@@
Table 1: Comparison showing the evolution of squared errors for models trained on different distributions; index N: $D_{F}=N(0,1)$, B $D_{F}=0.5N(-1,1)+0.5N(1,1)$ and U $D_{F}=U(-5,5)$. We show error rates for models prompted without and with the natural ordering on the prompts [sorted], for the large model size. $D^{t}_{i}=U(-1,1)$ and $D^{t}_{F}=N(0,\sigma)$
@@ 4.3 Error analysis, sigmoid approximations and boundary values
Our models’ performance depends on how often it has seen examples “similar” to the target function value it is trying to predict. At first, we thought this was due to the choice of coefficients in the target function $f(x)=ax+b$. However, experimentally, we verified that this is really due just to the values in the sequences it has seen. Extreme examples for $D_{F}=N(0,1)$ with tests in $[100,101]$ are in figure 2. In Appendix C we illustrate quantitatively intervals $I$ within which models have seen a large majority of values of sequences given a different training regime. Given a pretraining with over a billion examples, models will have seen prompts for functions with outside of $I$, just not many of them. As the models are tested with $D^{t}_{F}=N(0,\sigma)$ and so required to predict $\hat{f}(x)$ for $f(x)\not\in[-2,2]$, all the models do less and less well; Figure 5 in the Appendix shows similar behavior for models trained on uniform distributions.
Figure 2: Plots for model 12L8AH, trained on $D_{I}=D_{F}=N(0,1)$ for $f(x)=x$ for high values (left) of $x$ and $f(x)=10x$ for normal (middle) then for low values of $x$ (right)
@@
This motivated us to investigate errors our models made for target functions $f(x)\not\in[-2,2]$– i.e. the values of $\hat{f}(x)$ outside the interval that includes the vast majority they have seen. Our models exhibit problematic behavior of 2 kinds. Even our best models, for $f(x)\not\in[-2,2]$ but reasonably close, say in $[-9,9]$, predict $\hat{f}(x)$ to a sigmoid-like function with correct estimates for the target function within a certain interval. Consider the middle plot for $f(x)=10x$ in Figure 2. The plot shows that the model’s prediction $\hat{f}(x)$ diverges dramatically from $f(x)$ outside of a certain interval, but the rightmost plot shows that it has approximated well within that interval. Appendix D contains a graph over length of the prompt showing that it has learned something with ICL.
@@
For equations $f(x)$ sampled outside $N(0,1)$ (for example $f(x)=30x+30$ and $D^{t}_{I}=N(0,1)$, however, the results are catastrophic and similar to those in the first plot of Figure 2. Figure 4 in the Appendix shows that the model doesn’t converge to any stable prediction with ICL.
@@
This behavior across a wide range of models. For example with $D_{F}=D_{I}=U(-5,5)$, consider again as an illustrative example the target function, $f(x)=9x$ for our largest trained model. The model approximates $f(x)$ well within a certain range $[-B,B]$, but it predicts $\hat{f}(x)$ to be a constant function for $x$ such that $\hat{f}(x)\not\in[-B,B]$ within a certain range (See Figure 5 and discussion in Appendix C). We call values $-B,B$ boundary values. By training on uniform distributions, we can determine the boundary values exactly; e.g, for $U(-5,5)$ $B=5\times 5+5$. These are the biggest and smallest values the model could have seen during training. If such a model hasn’t seen a value above B or below -B, it won’t infer one. Different models trained on different uniform distributions give different boundary values (see below).
@@
All our models trained on $U(-5,5)$ estimate the target function more or less well for $x$ with $f(x)\in[-30,30]$ ; but once we are outside $[-B,B]$, the estimations become constant functions or chaotic. Figure 5 with equation $f(x)=40x+40$—illustrates this chaotic behavior as does the leftmost plot of Figure 2 for function $f(x)=x$ with large number inputs.
To summarize:
Observation 4
(i) For all models $M$ and for values $B<f(v)<B+\alpha$, where $\alpha$ is a constant determined by $M$, $fh_{M}(v)\approx B$, and for $-B-\alpha<f(v)<-B,fh_{M}(v)\approx-B$. (ii) For functions and data samples when the values of $f(x)$ in the prompt sequence are such that $f(x)>B+\alpha$ or $<-B-\alpha$, the model assigns $\hat{f}(v)$ random values for $f(v)$ far away from $B$ (i.e $>B+\alpha$ or $<-B-\alpha$.
@@
Constraints from boundary values hold for all transformer models tested (for plots see Appendix D and Figure 6) and for attention only models (See Appendix D, Figure 8). However, due to the parameter $\alpha$, larger models trained on the same distribution and the same number of data will ICL ${\cal L}$ functions over a slightly larger number of intermediate values than smaller models, as Figure 1 suggests. Figure 7 in the appendix shows plots for the predictions of two models (12L8AH, and 6L4AH) for $D_{F},D_{I}=N(0,1)$ for target $f(x)=10x$. The larger model has boundary values $\approx$ -13.7, 13.7, the smaller one boundary values $\approx$ -12, 12.
Giannou et al. (2024) also noted something like boundary values with their linear transformer architecture but they do not accord them much importance. They also investigated out of distribution behavior but only on $D_{I}\neq D^{t}_{I}$ (covariate shifts in Zhang et al. (2024)) (not shifts from $D_{F}$). They found that after 4 layers transformer model performance did not perform. We found that larger models did improve performance, but when we set $D_{I}\neq D^{t}_{I}$, we got bad results when the function’s values on those points were outside what we call boundary values, something which held for all models.
Zhang et al. (2024)’s covariate shift is also different from our experiments. They shift the prompt distribution but not that of the query. When we take a distribution over input points in train $D_{I}$ and set $D^{t}_{I}\neq D_{I}$, our shift is not the same; we shift both prompt and query distributions. With covariate shifts we found that the choice of points is important and model performance degrades considerably when the values of the functions on the chosen points lie beyond what we call boundary values, which Zhang et al. (2024) do not. As far as we know we are the first to take boundary values and their dependence on model parameters as key indications of what is actually going on in ICL.